OpenStudy (anonymous):

Last week Brandon and Marcee paid a total of 17$ to go to the movies. This week six of their friends are joining them. How much can they expect to pay altogether? $34 $51 $68 $102

OpenStudy (amistre64):

if we are to assume that 17 is just the price of admission and not any extra luxuriies; then 17/2 is the price for 1 and there are now 8 of them

OpenStudy (amistre64):

17*8 ---- = 17*4 = 40+28 = ? 2

OpenStudy (anonymous):

$68

OpenStudy (amistre64):

id go with 68 as well :)

OpenStudy (anonymous):

I'd say the group is 4 times larger...so 4(17) = 68. :)